solve this equation y2=144 and round to the nearest tenth
Answer by JBarnum(2146)

its the square root of 144 which is 12, y=12
because the square root of y^2=y
a good thing to practice and memorize are all the squares from 2^2 up to 12^2
2^2=4
3^2=9
4^2=16
5^2=25
6^2=36
7^2=49
8^2=64
9^2=81
10^2=100
11^2=121
12^2=144
I say this as these Squares are most commonly Used in school books so if u already know the relation then these type of questions become a 2sec question on a test.